# TableEval: A Real-World Benchmark for Complex, Multilingual, and Multi-Structured Table Question Answering

### v2

@@ A.2 Multi-answer Consistency Checking
A.3 Structured Answer Extraction
Appendix B Experimental Setup
We evaluate 19 models, ranging from 7B to 671B parameters, to assess their performance on TableEval. The evaluated models include open-source and closed-source models to comprehensively compare different model architectures and capabilities. For open-source models, we evaluate Qwen2.5-Instruct series Qwen et al. (2025): Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct; Qwen2.5-Coder-Instruct-32B Hui et al. (2024); QwQ-32B-Preview; DeepSeek series DeepSeek-AI et al. (2024): DeepSeek-V2-Lite-Chat (16B), DeepSeek-Coder-V2-Lite-Instruct (16B), DeepSeek-V2.5-1210 (236B), DeepSeek-V3 (671B), DeepSeek-R1 (671B) DeepSeek-AI et al. (2025); glm-4-9b-Chat (9B) GLM et al. (2024); Llama-3 series Grattafiori et al. (2024): Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct. For closed-source models, we evaluate qwen-max-2024-09-19, claude-3.5-sonnet-20241022, GPT-4o-mini-2024-07-18, GPT-4o-2024-11-20 OpenAI et al. (2024), and o1-preview. We set the temperature to 0 and use a fixed seed (42) to make the model’s output more deterministic. For models that do not allow the configuration of parameters (e.g., temperature or seed), we use their default settings.
Appendix C Additional Experimental Analysis
C.1 Robustness of SEAT
C.2 Performance on Sub-tasks
It can be found in the information retrieval sub-tasks (simple, condition, and grouped lookups) that larger or more specialized models (particularly DeepSeek-R1, claude-3.5-sonnet, GPT-4o, and QwQ-32B-Preview) consistently achieve the highest performance. QwQ-32B-Preview and DeepSeek-R1 better handle more complex filtering (condition, grouped), suggesting that the enhanced reasoning capabilities help manage varied data partitions. A similar pattern is shown in numerical analysis (sorting, statistics, calculation), with claude-3.5-sonnet exceeding 95% on sorting and DeepSeek-R1 exceeding 86% on statistics and calculation. This performance indicates that large model sizes and specialized data or method-specific fine-tuning can improve arithmetic-driven tasks.
For reasoning, we focus on multi-hop questions requiring chained inference across table cells. DeepSeek-R1 achieves a top score of 87.91%, followed by closed-source models such as claude-3.5-sonnet and GPT-4o, highlighting the benefits of size and specialized training. Smaller open-source models (e.g., glm-4-9b-chat, DeepSeek-V2-Lite-Chat) fall behind, reflecting the complexity of multi-hop inference towards small-scale models.
Data analysis includes six sub-tasks (causal analysis, comparative analysis, trend analysis, anomaly detection, correlation judgment, and rejection). These tasks require higher-level inference and domain knowledge, and the results reveal significant performance spread across models. For example, Llama-3.3-70B-Instruct achieves high scores in causal and trend analysis, while QwQ-32B-Preview achieves one of the highest scores on anomaly detection. Conversely, o1-preview demonstrates strong correlation judgment (78.40%), and claude-3.5-sonnet outperforms all other models on rejection (92.05%), reflecting its robustness when queries demand domain knowledge or interpretation of numeric relationships. The difference in these data analysis tasks suggests that even leading LLMs specialize differently, implying opportunities for method-level customizations to boost performance on specific analytical tasks.
In the ellipsis/reference sub-task, claude-3.5-sonnet achieves the best (87.94%), followed closely by DeepSeek-V3. Large-scale open-source models (e.g., Llama-3.3-70B-Instruct, Qwen2.5-72B-Instruct) also obtain high scores, yet they still fall behind top closed-source systems, demonstrating the benefit of extensive multi-turn dialogue training data in proprietary setups.
@@ C.2 Performance on Sub-tasks
C.3 Impact of Table Format
@@
In sub-task evaluations, all models exhibit around a 2$\times$ improvement in merged cell detection on HTML tables. Additionally, in table size detection, all models show varying degrees of progress. These results indicate that structured table formats (HTML and LaTeX) significantly enhance the models’ ability to understand table structures, leading to improved TableQA performance.
C.4 Impact of Fine-tuning
@@
Our experimental results demonstrate that fine-tuning significantly enhances the TableQA capabilities of LLMs. Even small-scale LLMs can achieve comparable performance with GPT-4o-mini after fine-tuning. Specifically, the fine-tuned Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct achieved notable improvements, particularly in table structure understanding, where they even outperform GPT-4o-mini. These results suggest that fine-tuning with task-specific TableQA data can help models develop a more structured understanding of tables.
@@
Appendix D Prompts
@@
Figure 6: QA generation prompt in role-prompted strategy.
Figure 7: QA generation prompt in template-prompted strategy.
@@
2:     $\mathcal{T}\leftarrow\texttt{LocateTables}(\mathcal{L})$triangleright Identify all tables in the layout
@@
4:         $C_{\text{up}}\leftarrow\emptyset$triangleright Initialize upward context
5:         $C_{\text{down}}\leftarrow\emptyset$triangleright Initialize downward context $\triangleright$ Upward Context Search
6:         for each Window $w$ in TraverseUpward($T,\omega$do
7:              if IsHeader($w$then
@@
10:              else if IsNonTextContent($w$then
@@
13:              else if CharCountExceeded($C_{\text{up}},\kappa$then
@@
15:              else if WindowRangeExceeded($w$then
@@
19:         end $\triangleright$ Downward Context Search
20:         for each Window $w$ in TraverseDownward($T,\omega$do
21:              if IsHeader($w$then
@@
24:              else if IsNonTextContent($w$then
@@
27:              else if CharCountExceeded($C_{\text{down}},\kappa$then
@@
29:              else if WindowRangeExceeded($w$then
@@
36:     return $\mathcal{T}$triangleright Return tables with augmented context
@@
38:function $\mathcal{L}$)
[…]

### v3

@@ B.1 Robustness of SEAT
B.2 Performance on Sub-tasks
It can be found in the information retrieval sub-tasks (simple, condition, and grouped lookups) that larger or more specialized models (particularly DeepSeek-R1, Claude 3.5 sonnet, GPT-4o, and QwQ-32B-Preview) consistently achieve the highest performance. QwQ-32B-Preview and DeepSeek-R1 better handle more complex filtering (condition, grouped), suggesting that the enhanced reasoning capabilities help manage varied data partitions. A similar pattern is shown in numerical analysis (sorting, statistics, calculation), with Claude 3.5 sonnet exceeding 95% on sorting and DeepSeek-R1 exceeding 86% on statistics and calculation. This performance indicates that large model sizes and specialized data or method-specific fine-tuning can improve arithmetic-driven tasks.
For reasoning, we focus on multi-hop questions requiring chained inference across table cells. DeepSeek-R1 achieves a top score of 87.91%, followed by closed-source models such as Claude 3.5 sonnet and GPT-4o, highlighting the benefits of size and specialized training. Smaller open-source models (e.g., glm-4-9b-chat, DeepSeek-V2-Lite-Chat) fall behind, reflecting the complexity of multi-hop inference towards small-scale models.
Data analysis includes six sub-tasks (causal analysis, comparative analysis, trend analysis, anomaly detection, correlation judgment, and rejection). These tasks require higher-level inference and domain knowledge, and the results reveal significant performance spread across models. For example, Llama-3.3-70B-Instruct achieves high scores in causal and trend analysis, while QwQ-32B-Preview achieves one of the highest scores on anomaly detection. Conversely, o1-preview demonstrates strong correlation judgment (78.40%), and Claude 3.5 sonnet outperforms all other models on rejection (92.05%), reflecting its robustness when queries demand domain knowledge or interpretation of numeric relationships. The difference in these data analysis tasks suggests that even leading LLMs specialize differently, implying opportunities for method-level customizations to boost performance on specific analytical tasks.
In the ellipsis/reference sub-task, Claude 3.5 sonnet achieves the best (87.94%), followed closely by DeepSeek-V3. Large-scale open-source models (e.g., Llama-3.3-70B-Instruct, Qwen2.5-72B-Instruct) also obtain high scores, yet they still fall behind top closed-source systems, demonstrating the benefit of extensive multi-turn dialogue training data in proprietary setups.
@@ B.2 Performance on Sub-tasks
B.3 Impact of Table Format
@@
[…]
B.4 Impact of Fine-tuning
@@
Our experimental results demonstrate that fine-tuning significantly enhances the TableQA capabilities of LLMs. Even small-scale LLMs can achieve comparable performance with GPT-4o-mini after fine-tuning. Specifically, the fine-tuned Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct achieved notable improvements, particularly in table structure understanding, where they even outperform GPT-4o-mini. These results suggest that fine-tuning with task-specific data can help models develop a more structured understanding of tables.
@@
Appendix C Prompts
Figure 5: QA generation prompt in role-prompted strategy.
@@ Appendix C Prompts
Figure 9: QA generation prompt in template-prompted strategy.
@@
2:  $\mathcal{T}\leftarrow\texttt{LocateTables}(\mathcal{L})$ $\triangleright$ Identify all tables in the layout
@@
4:   $C_{\text{up}}\leftarrow\emptyset$ $\triangleright$ Initialize upward context
5:   $C_{\text{down}}\leftarrow\emptyset$ $\triangleright$ Initialize downward context $\triangleright$ Upward Context Search
6:   for each Window $w$ in TraverseUpward($T,\omega$) do
7:     if IsHeader($w$) then
@@
10:     else if IsNonTextContent($w$) then
@@
13:     else if CharCountExceeded($C_{\text{up}},\kappa$) then
@@
15:     else if WindowRangeExceeded($w$) then
@@
[…]
20:   for each Window $w$ in TraverseDownward($T,\omega$) do
21:     if IsHeader($w$) then
@@
24:     else if IsNonTextContent($w$) then
@@
27:     else if CharCountExceeded($C_{\text{down}},\kappa$) then
@@
29:     else if WindowRangeExceeded($w$) then
@@
36:  return $\mathcal{T}$ $\triangleright$ Return tables with augmented context
@@
38:function LocateTables($\mathcal{L}$)
39:  return ElementsOfType($\mathcal{L},\text{``Table''}$)